OpenStudy (anonymous):

suppose that the nation demand and price for a certain type of energy-efficient exhaust fan are related by p=490-5/7(q), where p is the price in dollars and q is the demand in thousands. if the demand is for 161000 exhaust fans, than the price is?

OpenStudy (anonymous):

the equation that the problem is simply an equation of a line... if we simply plug-in the given demand q=161000, you can solve the equivalent price p... :)

OpenStudy (anonymous):

can you do it?

OpenStudy (anonymous):

−114510

OpenStudy (anonymous):

this is what i get but it is not right

OpenStudy (anonymous):

we know that the price p can't be a negative value so we have to determine the boundaries of our equation based on what is needed...

OpenStudy (anonymous):

I've tried everything and can't seem to get it

OpenStudy (anonymous):

if price p=0, we can determine demand q as \[0=490-\frac{5}{7}q\]\[-490=\cancel{-490}+\cancel{490}-\frac{5}{7}q\]\[-490=-\frac{5}{7}q\]\[-\frac{7}{5}\times-490=\cancel{-\frac{5}{7}}q\times\cancel{-\frac{7}{5}}\]\[q=\frac{7\times490}{5}=686\]

OpenStudy (anonymous):

therefore we cannot have a demand q more than 686...

OpenStudy (anonymous):

oh I see

OpenStudy (anonymous):

since at this point our price p=0

OpenStudy (anonymous):

but if we don't have a demand q=0, our price p=490

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did i explain it clear? you get the point? :)

OpenStudy (anonymous):

yes I get it now thank you very much

OpenStudy (anonymous):

no problem, always glad to help if it is needed... :)

OpenStudy (anonymous):

:)